The Ryde and Rowe Inc. had the following account balances as of January 1 Inventory Inventory Inventory Overhead Direct Materials Work in Process Finished Goods Manufacturing $ 8,700 76,500 53,000 0 During the month of January, all of the following occurred 1. Direct labor costs were $48,000 for 1,800 hours worked 2 Direct materials costing $29,000 and indirect materials costing $5,100 were purchased 3. Sales commissions of $17,000 were earned by the sales force 4. $22,000 worth of direct materials were used in production. 5. Advertising costs of $6,300 were incurred 6. Factory supervisors earned salaries of $11,935 7. Indirect labor costs for the month were $3,000 8. Monthly depreciation on factory equipment was $4,500 9. Utilities expense of $6,899 was incurred in the factory 10. Equipment with manufacturing costs of $69,000 were transferred to finished goods 11. Monthly insurance costs for the factory were $4,200 12. $5,000 in property taxes on the factory were incurred and paid 13. Equipment with manufacturing costs of $94,331 were sold for $171,511

Instructions a. Assume If Ryde and Rowe assigns manufacturing overhead of $34,400, what will be the balances in the Direct Materials, Work in Process, and Finished Goods Inventory accounts at the end of January? Direct materials inventory Work in process inventory Finished goods inventory b. As of January 31, what will be the balance in the Manufacturing Overhead account after the application of the overhead from part a? Manufacturing overhead

c. What was Ryde and Rowes operating income for January? Operating income
Answer #1

a) Ending balance of Direct material inventory = 8700+29000-22000 = 15700

Ending balance of work in process = 76500+22000+48000+34400-69000 = $111900

Ending balance of finished goods = 53000+69000-94331 = $27669

b) Balance in manufacturing overhead :

Actual manufacturing overhead = 11935+3000+4500+6899+4200+5000 = 35534

Assign manufacturing overhead = 34400

Ending balance of manufacturing overhead = 35534-34400 = 1134 Debit balance

c) Operating income = 171511-94331-17000-6300 = 53880